OpenStudy (anonymous):

A sphere has a radius of 4x + 4. Which polynomial in standard form best describes the total surface area of the sphere? Use the formula S = 4r^2 for the surface area of a sphere.

OpenStudy (anonymous):

Please Help, I will give you a medal!

OpenStudy (stamp):

\[r=4x+4=4(x+1)\]\[S=4r^2\]Plug in r and simplify.

OpenStudy (anonymous):

4x^2+4 ? That doesn't seem right

OpenStudy (stamp):

\[r=(4x+4)\]\[S=4r^2\]\[S=4(4x+4)^2\]

OpenStudy (stamp):

Pay attention to your properties of substitution, and order of operations.

OpenStudy (anonymous):

64x^2+64 still doesn't seem right

OpenStudy (anonymous):

@stamp you have the formula for the surface area wrong. SA = (4)(pi)(r^2) Substituting, SA = (4)(pi)(4x + 4)^2

OpenStudy (stamp):

@tcarroll010 I am using the equation given to me by the op.

OpenStudy (anonymous):

What is the "op"? Whatever the "op" is, it/he/she is wrong. You left out : \[\Huge{\pi}\]

OpenStudy (anonymous):

64(pi)x^2 + 128(pi)x+ 64(pi)

OpenStudy (stamp):

@tcarroll010 op = original poster, the person who created the thread.

OpenStudy (anonymous):

Then correct the formula. Don't use an incorrect formula. It will lead to incorrect results.

OpenStudy (anonymous):

Stamp, aren't you the one who replied the first equation?

OpenStudy (stamp):

@tcarroll010 Well that is why we have you, friend. And I did not realize it was the incorrect formula, I normally reserve for OP the job of accurately providing the given resources required to solve a problem. Anyway, ya'll got this handled.

OpenStudy (anonymous):

So, if the formula: SA = (4)(pi)(r^2) is used for the surface area of a sphere, we can substitute: SA = (4)(pi)(4x + 4)^2 and then simplify.

OpenStudy (anonymous):

I got it now

OpenStudy (anonymous):

That's a good thing. Nice working with you, good luck in all of your studies and thx for the recognition! @ballerina789

OpenStudy (anonymous):

thanks. Your welcome